News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 21 March 2003 7.00am EST

During the last week Chandra completed the observing schedule as planned.

Last week we started collecting data from the HRC anti-coincidence shield at a lower than normal HV level during times when the HRC is not the observing instrument. These measurements are being performed as part of a study of the operations and flight software changes that would be necessary in order to use the HRC anti-coincidence shield to supplement or replace the EPHIN in our autonomous radiation protection system.

Of note last week was the Chandra Cycle 5 proposal deadline at 7pm EST on March 14. A total of 788 proposals were received and the submission process proceeded smoothly.

The observing schedule for next week is shown below and is largely devoted to a raster of over 60 observations of an NOAO field.

2MASS150706-122516     ACIS-S       Mar 23
HD134946               ACIS-I
SGR 1627-41            ACIS-S
SN 2003BG              ACIS-S       Mar 24
Radiation Belts                     Mar 25
IRAS04392-0123         ACIS-I
B2 0920+31B            ACIS-S
FIRST J0941+4121       ACIS-S
NOAO Field (23 obs)    ACIS-I       Mar 26
Radiation Belts                     Mar 27
NOAO Field (30 obs)    ACIS-I       Mar 28
Radiation Belts                     Mar 30
NOAO Field (9 obs)     ACIS-I


All spacecraft subsystems, with the exception of MUPS thruster A-1 as noted above, continued to operate nominally.

Return to Status Reports